# Unsettling the Hegemony of Intention: Agonistic Image Generation

### v2

@@ A.1. Participant Backgrounds
Table 2. Participant IDs with their selected prompt, category, and the order in which they encountered each of the non-Baseline interfaces.
@@ B.1. Design Iteration for Agonistic
In the first iteration of Agonistic, we generated different set of suggestions for each subject in the user’s prompt (an example with only one subject is shown in 7). Finding that generating suggestions for multiple subjects was too costly and slow for interview purposes, we stopped organizing suggestions by subject and instead only generated suggestions for the main subject of the prompt, as shown in 1.
@@
Figure 7. First iteration of Agonistic interface
@@ Appendix C Inter-Rater Reliability Methodology
We calculated inter-rater reliability (IRR) on value codes. the two first co-authors first independently coded a subset of 3 participant interviews using the coding ontology described in §4.3. Since some codes were not mutually exclusive with each other, we calculated a separate Cohen’s kappa score for each code, with the following results: 0.64 (Direct), 0.00 (Reminder), 0.65 (Expansion), 1.00 (Challenge), 0.66 (Realism), 0.83 (Familiarity), 0.60 (Diversity), 0.38 (Aesthetics) We then took an average of Cohen’s kappa scores across value codes, weighted by the frequency of each value code, to calculate our final IRR of 0.67.
We noticed that the Cohen’s kappa score for several codes might be unreliable to the low frequency of certain codes in the subset of interviews used to calculate IRR (for instance, there were only 3 instances of Reminder  and 1 instance of Challenge ). This issue was especially severe for intent codes, due to the extremely high proportion of Direct codes compared to other intents. For this reason, two first co-authors then reviewed all non-Directintent codes together to establish greater agreement. We note that while this approach reduced the chance of false positives for non-Directintent codes, it did not reduce the chance of false negatives. Therefore, the amount of reflection captured by our coding in this study is a conservative estimate of the amount of reflection that may have occurred.
@@ D.3. Change in design statement
Participants write an “design statement” — one or two sentences describing the most important aspects of the subject represented in the collage – after producing an initial collage with Baseline. The design statements provided both a more standardized and a richer way to record the progression of participant assumptions over different interfaces beyond the data collected by the self-reported rankings. After iterating through each following interface, participants copied over their previous design statement and made modifications to reflect the changes to their collage. They could opt to make no changes even if images were replaced, should they feel that the previous design statement satisfactorily described the new collage. The design statement reflects a different dimension of reflection than either the change in mental image or how images are added: for example, some participants did not report a substantive changes in mental picture of the subject but still modified the design statement to reflect a change in how they viewed their collage, like by emphasizing some aspects of the collage; some participants also replaced images in the collage but did not change the design statement because they felt that it applied just as well (the changes fell within the scope of the previous design statement). The design statement uniquely captures the self-perception of the participant as a collage designer synthesizing their choices.
@@
Table 3. Mean change in participant design statement after using each interface, as measured by Levenshtein and embedding-based distance. “Scaled” indicates 0-1 min-max scaling, where the smallest change is mapped to 0 and the highest is mapped to 1.
@@ D.5. Table Form for Results
Interface Rethink Appropriateness Control Interestingness
Baseline $2.03\pm 1.00$ $3.31\pm 0.91$ $3.03\pm 1.22$
Diverse  $1.48\pm 0.93$ $3.07\pm 1.20$ $2.59\pm 1.16$
Reformulative $1.93\pm 1.05$ $3.93\pm 0.91$ $3.79\pm 1.21$ $4.00\pm 1.11$
Agonistic $2.97\pm 1.19$ $3.83\pm 1.15$ $3.90\pm 0.99$ $4.28\pm 0.91$
@@
Interface Intents
Direct Reminder Expansion Challenge
Baseline 0.98 $0.00$ $0.02$ $0.00$
Diverse  0.95 $0.00$ $0.05$ $0.00$
Reformulative 0.90 $0.04$ $0.06$ $0.00$
Agonistic 0.67 $0.11$ $0.19$ $0.04$
@@
D.6. Quantitative Coding Analysis
A notable difference between the quantitative analysis of intent code occurrences and 5.1 is the finding that Diverse induces greater degrees of ‘Reminder’ and ‘Expansion’ of intent than Baseline ($+0.02$ and $+0.04$, respectively), in contrast to self-reported findings that Baseline induces greater reflection than Diverse. These differences may be explained by the fact that coding for added images excludes some forms of reflection, such as reflection that does not explicitly correlate to a particular image or reflection caused by images that participants ultimately reject, as further discussed in §6. For this reason, the methodology of this section may more accurately capture the unique reflection due to each interface because it gives less weight to reflection from external causes like performing the task itself. The ‘hierarchy’ of interfaces suggested by our findings in this section is thus $\left[\textsf{{Agonistic}}~{}\gg\textsf{{Reformulative}}~{}>\{\textsf{{Baseline}},\textsf{{Diverse}}\}\right]$.
D.7. Additional Qualitative Themes
Agonistic can feel like a double-edged sword, treading the line between appropriate and inappropriate. The results in §6.1 suggest that appropriateness is correlated with higher rethinking. However, by design, Agonistic often presents users with interpretations that are perceived initially by users as unrelated, inappropriate, or even offensive. For many participants who underwent reflection, these interpretations eventually seemed appropriate, explaining the high appropriateness rating of Agonistic. However, for other participants, these interpretations remained inappropriate (even after reflection). Several participants commented on the dually appropriate and inappropriate character of Agonistic. P4 suggested that “it [Agonistic] was giving me the most appropriate, but it was also giving me the most inappropriate [outputs] at the same time… [it] would sometimes hit a home run, sometimes wildly off track.’P16 said that Agonistic was “good good and bad bad […] I like the specificity [of Agonistic]. I just think the specificity was kind of whacked out in a way where it just wasn’t super on the bar of what I was thinking about”. Together, these suggest that the experience of using Agonistic can vary widely, and that perhaps designing for reflection via agonism means that interface outputs will not seem appropriate to many users by its own nature.
Appendix E Extended Discussion
E.1. The Many Heads of Reflection
We find it important to note that people find reflection across all the interfaces, as demonstrated by Table 1; although the overall occurrence of reflection does vary substantively (as discussed in §5), reflection is still a very subjective, personal experience. Indeed, while Baseline was uninspiring for many — “I typed what I thought and I got what I thought I would getP21 – it was a highly reflective experience for others — “I have a lot of other potential images in my mind of it now, and I don’t know what is […] actually trueP28. We find that reflection generally seems to emerge from an initially unexpected but eventually rationalizable encounter with the interface (either the images themselves or contextualizing content). This reflects a host of work on reflection and related cognitive activities emphasizing encounter with alterity or the Other (Ye et al., 2024; Halbert and Nathan, 2015; Levinas, 1969; Lacan, 2006, inter alia). Although reflection often occurred in planned ways (i.e., as a result of design elements we had introduced with the intention of inducing reflection), it also often emerged in spontaneous but nonetheless personally significant ways. This emphasizes that image generation interfaces may benefit from more adaptive design, but more importantly that users will always reflect in ways that interface designers cannot completely control.

### v3

@@ A.1. Participant Backgrounds
Table 2. Participant IDs with their selected prompt, category, and assigned order for non-Baseline interfaces.
@@ B.1. Design Iteration for Agonistic
In the first iteration of Agonistic, we generated different set of suggestions for each subject in the user’s prompt (an example with only one subject is shown in 8). Finding that generating suggestions for multiple subjects was too costly and slow for interview purposes, we stopped organizing suggestions by subject and instead only generated suggestions for the main subject of the prompt, as shown in 1.
@@
Figure 8. First iteration of Agonistic interface.
@@ Appendix C Inter-Rater Reliability Methodology
We calculated inter-rater reliability (IRR) on value codes in the following way. First, the two first co-authors independently coded a subset of 3 participant interviews using the coding ontology described in §4.3. Since the codes were not mutually with each other, we calculated a separate Cohen’s kappa score for each code, with the following results: 0.64 (Direct), 0.00 (Reminder), 0.65 (Expansion), 1.00 (Challenge), 0.66 (Realism), 0.83 (Familiarity), 0.60 (Diversity), 0.38 (Aesthetics). We then took an average of Cohen’s kappa scores across value codes, weighted by the frequency of each value code, to calculate our final IRR of 0.67.
We noticed that the Cohen’s kappa score for intent codes was unusually low due to the high frequency of Direct codes compared to other intent codes in the subset of interviews used to calculate IRR (for instance, there were only 3 instances of Reminder and 1 instance of Challenge in the subset) (Feinstein and Chicchetti, 1990). For this reason, we did not include intent codes in our IRR calculations. Instead, the two first co-authors reviewed all non-Direct intent codes together to establish agreement. We note that while this approach reduced the chance of false positives for non-Direct intent codes, it did not reduce the chance of false negatives. Therefore, the amount of reflection captured by our coding in this study is a conservative estimate of the amount of reflection that may have occurred.
@@ D.3. Change in Design Statement
Participants write an “design statement” (one or two sentences describing the most important aspects of the subject represented in the collage) after producing an initial collage with Baseline. The design statements provided both a more standardized and a richer way to record the progression of participant assumptions over different interfaces beyond the data collected by the self-reported rankings. After iterating through each following interface, participants copied over their previous design statement and made modifications to reflect the changes to their collage. They could opt to make no changes even if images were replaced, should they feel that the previous design statement satisfactorily described the new collage. The design statement reflects a different dimension of reflection than either the change in mental image or how images are added: for example, some participants did not report a substantive changes in mental picture of the subject but still modified the design statement to reflect a change in how they viewed their collage, like by emphasizing some aspects of the collage; some participants also replaced images in the collage but did not change the design statement because they felt that it applied just as well (the changes fell within the scope of the previous design statement). The design statement uniquely captures the self-perception of the participant as a collage designer synthesizing their choices.
@@
Table 3. Mean change in participant design statement after using each interface, as measured by Levenshtein and embedding-based distance. “Scaled” indicates 0-1 min-max scaling, where the smallest change is mapped to 0 and the highest to 1.
@@ D.5. Table Form for Results
Interface Rethink Appropriateness Control Interestingness
Baseline $2.03\pm 1.00$ $3.31\pm 0.91$ $3.03\pm 1.22$
Diverse  $1.48\pm 0.93$ $3.07\pm 1.20$ $2.59\pm 1.16$
Reformulative $1.93\pm 1.05$ $3.93\pm 0.91$ $3.79\pm 1.21$ $4.00\pm 1.11$
Agonistic $2.97\pm 1.19$ $3.83\pm 1.15$ $3.90\pm 0.99$ $4.28\pm 0.91$
@@
Interface Intents
Direct Reminder Expansion Challenge
Baseline 0.98 $0.00$ $0.02$ $0.00$
Diverse  0.95 $0.00$ $0.05$ $0.00$
Reformulative 0.90 $0.04$ $0.06$ $0.00$
Agonistic 0.67 $0.11$ $0.19$ $0.04$